### v1

# Evaluating the Performance and Robustness of LLMs in Materials Science Q&A and Property Predictions

Abstract
Large Language Models (LLMs) have the potential to revolutionize scientific research, yet their robustness and reliability in domain-specific applications remain insufficiently explored. This study conducts a comprehensive evaluation and robustness analysis of LLMs within the field of materials science, focusing on domain-specific question answering and materials property prediction. Three distinct datasets are used in this study: 1) a set of multiple-choice questions from undergraduate-level materials science courses, 2) a dataset including various steel compositions and yield strengths, and 3) a band gap dataset, containing textual descriptions of material crystal structures and band gap values. The performance of LLMs is assessed using various prompting strategies, including zero-shot chain-of-thought, expert prompting, and few-shot in-context learning. The robustness of these models is tested against various forms of ‘noise’, ranging from realistic disturbances to intentionally adversarial manipulations, to evaluate their resilience and reliability under real-world conditions. Additionally, the study uncovers unique phenomena of LLMs during predictive tasks, such as mode collapse behavior when the proximity of prompt examples is altered and performance enhancement from train/test mismatch. The findings aim to provide informed skepticism for the broad use of LLMs in materials science and to inspire advancements that enhance their robustness and reliability for practical applications.
Introduction
Large Language Models (LLMs) represent a significant advancement in the field of artificial intelligence and have been rapidly adopted for application in various scientific disciplines [1]. With their ability to process and generate natural language, LLMs are potent tools for tasks like information retrieval, question and answering (Q&A), and property predictions [2, 3, 4]. Similar to traditional ML models, LLMs can require extensive data processing, large volumes of data, and massive compute resources to train [5]. Despite these limitations, pretrained LLMs can be adapted to new tasks with few-shot examples via in-context learning (ICL), making them both cost-effective and rapid to deploy [6, 7]. In the context of materials science, where data acquisition can often be costly and time-consuming, leveraging ICL enables LLMs to efficiently prototype and generate predictive insights even in low-data settings [8, 9, 10]. Recent work has demonstrated that LLMs are capable of domain-specific Q&A [11, 12, 13], materials property predictions [14, 15, 16], and information extraction from complex datasets [4, 17]. These studies demonstrate LLMs’ potential to serve as flexible and powerful analytical tools.
However, the robustness of LLMs is crucial in real-world scenarios, but has not been sufficiently investigated in the literature. One aspect of the robustness of LLMs is their sensitivity to prompt changes either due to innocuous or adversarial reasons [18, 19]. Variations in how a query or instruction is formulated may cause a response to factually change [18]. As an example, 0.1 nm and 1 Angstrom are equivalent but switching them in a prompt could result in different LLM predictions for the same task. Alternatively, the response of the LLM can be deliberately altered through intentional misinformation or misleading inputs [19]. These attributes are not only theoretical concerns but are critical for the reliable usage of LLMs as they become integrated into the materials science research and development pipeline. Given that LLMs generate outputs with indifference to truth [20], thoroughly probing LLM prompt sensitivity would allow us to critically evaluate model performance in practical situations; providing informed skepticism for the broad use of LLMs in materials science.
In this work, we conducted a holistic robustness analysis of commercial and open-source LLMs for materials science. We systematically investigated how textual perturbations like text ordering within a prompt, and common unintentional and intentional prompt variations influence the inference of LLMs. Three distinct datasets of domain-specific Q&A and materials property prediction were selected and first benchmarked via prompt engineering to identify baseline and optimal performance boundaries. We then investigated the impact of different types of textual perturbations, from realistic to adversarial perturbations, on the performance of various-sized LLMs in the materials science Q&A. We then used property predictions from the matbench_steels dataset to show that even without fine-tuning, pretrained LLMs demonstrated enhanced predictive ability through few-shot ICL when presented with similar examples to the prediction task. Conversely, when provided with dissimilar examples during few-shot ICL, mode collapse behavior was observed, where the model generated identical outputs despite varying inputs. Counterintuitively, for the fine-tuned LLM examined on band gap predictions, supposedly adversarial perturbations like shuffling or randomization enhanced the model’s predictive capability. Interestingly, this enables the fine-tuned model to function effectively even with significantly ablated prompts. This train/test mismatch behavior, absent in traditional ML models, highlights a potential direction for distilling LLM-based predictive models.
Results
An overview of the framework is illustrated in Figure 1. For the performance evaluation and robustness analysis of materials science Q&A, we use the MSE-MCQs dataset, which has 113 multiple-choice questions from a first-year introductory materials science and engineering course at the University of Toronto. The questions are meant to test students’ understanding of materials science knowledge, including material mechanics, thermodynamics, crystal structures, materials properties, etc. The questions are categorized into easy (number of questions, n=39), medium (n=40), and hard levels (n=34), based on the conceptual difficulty, calculation complexity, and the depth of reasoning required. For the performance evaluation of property prediction, we use matbench_steels, a subcategory of the Matbench test set originally proposed for benchmarking traditional machine learning (ML) models for materials property predictions [21]. The matbench_steels dataset has 312 pairs of material compositions (as chemical formula) and yield strength (in MPa). For the robustness analysis of property prediction, we use a band gap dataset, which comprises 10,047 descriptions of material crystal structures generated via Robocrystallographer, along with their corresponding band gap values from the Materials Project database [14].
Figure 1: Schematic representation of the experiment design for performance evaluation and robustness analysis of LLMs. Yellow highlights the testing conducted in Q&A settings. Blue highlights the testing conducted in property prediction settings. Green represents the tests associated with performance evaluation. Red represents tests related to degradation and robustness analysis.
Performance Evaluation of Materials Science Q&A
The results of the performance evaluation of LLMs on the MSE-MCQs dataset are shown in Figure 2. For each model and setting, three trials were conducted, and the error bars represent the inherent non-determinism in LLMs even at the lowest temperature settings. This non-determinism may affect the reliability and reproducibility of the models [22] and is likely a result of the stochastic sampling during text generation [23, 24, 25]. The result shows that gpt-4-0613 scored the highest in all categories, followed by gpt-3.5-turbo-0613 and llama2-70b. The remaining models scored around or slightly below the baseline score of 0.246, which is equivalent to random guessing for the provided MCQs.
Figure 2: LLM performance evaluation and prompt engineering enhancement in materials science Q&A using MSE-MCQs dataset. On each bar, the lighter color represents the performance of those models when introduced with the expert prompt. Error bars represent the standard deviation due to LLMs’ non-determinism.
Upon implementing the expert prompt, we observed a consistent improvement in model performance across almost all the models and question types. For gpt-3.5-turbo-0613, the significant variation in the data indicated by the large error bar prevents a definitive conclusion about the expert prompt improving performance on easy and medium-level questions. However, the accuracy of hard-level questions nearly doubled.
We investigated why the smaller Llama models (llama-13b and llama-7b) scored lower than the baseline without the expert prompt. Beyond the limited skill levels, these models even failed to understand the intent when instructions were not provided. Instead of answering, they often attempted to “complete” the questions rather than answering them. However, once the expert prompt was implemented, these smaller models could follow the instructions and attempt to solve the questions, in which case the performance improved to around and sometimes above the baseline scores.
The performance trends observed when evaluating large language models on the MSE-MCQs dataset were expected. Specifically, larger models like gpt-4-0613 significantly outperformed smaller ones, with model performance generally improving across the board when an expert prompt was used, especially in handling more complex questions.
Performance Evaluation of Materials Property Prediction
Here we investigate LLM materials property predictions, utilizing the matbench_steels dataset. Figure 3 presents a comparative analysis between gpt-3.5-turbo-0613, employing few-shot ICL, and a traditional random forest regressor (RFR) model, in predicting the yield strength of steels. It is important to note that gpt-3.5-turbo-0613 was not subject to an explicit training phase for this task but processed the data as few-shot examples. In contrast, the RFR was trained explicitly using MAGPIE features from the same data points [26].
Figure 3: Prediction performance vs. number of training points of RFR and gpt-3.5-turbo-0613. The distributions indicate models’ sensitivity to the selection of training points, both existing in the LLM model with ICL and RFR model with traditional training process. In each violin, the center dashed line represents the median, and the upper and lower dashed lines represent the first and third quartile, respectively.
Figure 3 summarizes the distribution of errors for the two models as a function of the number of training points used. It can be observed that gpt-3.5-turbo-0613, utilizing its ICL capabilities, performs comparably to the trained RFR model. When given only 5 data points as few-shot examples, the median MAE of gpt-3.5-turbo-0613 begins at a slightly higher value of 249 MPa but improves significantly to 171 MPa when increased to 25 data points. Meanwhile, the RFR model, starting with a slightly lower median MAE of 231 MPa with a large error bar, shows a more modest improvement to 194 MPa at 25 training points. The declining MAE trend suggests that both models benefit from more data, yet the LLM’s rate of improvement is more significant.
This analysis reveals a notable adaptability of pretrained LLMs to new predictive challenges using ICL, particularly when data availability is constrained. Such adaptability makes LLMs potentially valuable in early-stage research or exploratory studies in materials science, where data may be scarce or costly to obtain. However, as the number of data points increases, most LLMs suffer from limited prompt windows, which make such applications computationally expensive or impossible, in which case fine-tuned LLMs and traditional machine learning models with dedicated training may be more effective.
To better understand how the selection of highly relevant training data can be used to enhance LLM performance we conducted a systematic study using nearest-neighbor-boosted ICL. The method developed here involves deliberately selecting training data points based on representational proximity to enhance the model’s performance. Figure 4 shows the prediction performance when LLMs are tasked with predicting yield strength using a) farthest neighbors, b) random neighbors, and c) nearest neighbors. For performance comparison, a k-nearest neighbors (KNN) model and RFR were also implemented.
Figure 4: Prediction performance of various LLMs under different training neighbor settings. Three panels compare the model performance across neighbor selection methods (left to right): Farthest Neighbors, Random Neighbors, and Nearest Neighbors. The TPOT-Mat performance is indicated by a horizontal grey line for benchmarking.
When trained with farthest neighbors, models generally exhibit a decrease in MAE as the number of neighbors increases. A notable exception to this trend is llama2-13b, which shows a non-monotonic decrease in MAE. The GPT models both outperformed the KNN and RFR models when fed with more than 15 points. This indicates that LLMs may have an inherent capability to extract valuable patterns even from data points that are less obviously related to the test set.
For the random neighbors training set, the LLMs’ performance consistently improves as the dataset size increases. This suggests that a randomly composed dataset offers a balanced and rich learning environment for these models, allowing for the development of more robust generalizations about the data. The GPT models outperform the RFR model as the data size increases, indicating that their more sophisticated architectures and larger training corpora enhance their ability to analyze and interpret complex data relationships more effectively. On the other hand, the smaller LLMs exhibit less stable performance with higher MAE values which exhibit non-monotonic variation with the number of neighbors. The random neighbors setting appears to challenge these models to a greater degree, potentially because their smaller scale and limited pretraining are less capable of handling the diverse input without fine-tuning or additional data processing.
The nearest neighbors represent the most relevant data points in the compositional space to the test points. The KNN model exhibits an increase in MAE with a growing number of neighbors, which is expected as additional (more distant) neighbors are less similar to the predicted sample. This suggests that if LLMs act purely on information provided without additional internal calculation, their performance should degrade as the number of neighbors increases. The Llama models do not perform as well as the KNN or RFR models. Contrastingly, as the data size increases, the larger LLMs show a consistent decrease in MAE and outperform the RFR model. While the gpt-4-0613 model with 20 nearest neighbors achieves an MAE of 89.6, it does not surpass the best-performing ML model, TPOT-Mat, which has an MAE of 79.9 on the matbench_steels dataset [27]. However, it is important to note that this is not a direct comparison, as TPOT-Mat employs a 5-fold nested cross-validation method on Matbench datasets [21], which utilizes all available data and is likely to result in better performance. This highlights the promise of LLMs in data-lean materials property prediction tasks without the need for feature engineering, particularly given their general-purpose design and lack of task-specific fine-tuning.
To investigate the model’s predictive behaviors under these different settings, we analyzed the parity plots of the gpt-4-0613’s predictions when utilizing 20 neighboring data points, as shown in Figure 5. Alongside these plots, the figure also includes histograms of the top five most frequently predicted yield strength values, to ensure the model is not merely guessing a few commonly present values. This behavior is known as “mode collapse”, whereby a generative model can favor a certain output due to overfitting to its pretraining data or lack of generalization capability [28]. Understanding mode collapse is crucial for evaluating the robustness of LLMs because it directly impacts the model’s reliability and utility in practical applications. By identifying the mode collapse behavior, one can evaluate the quality of those predictions and potentially improve the performance.
Figure 5: Parity plots (top) and associated mode collapse point histograms (bottom) of gpt-4-0613 under different training neighbor settings: Farthest Neighbors, Random Neighbors, and Nearest Neighbors, from left to right.
In the farthest neighbors setting, the discrete nature of the predictions, as evidenced by the observation of horizontal lines in the figure, suggests that the model is unable to capture the relationship between the composition and the yield strength effectively. The histogram indicates a strong mode collapse behavior, with the model frequently predicting a handful of values. This suggests that the model may be defaulting to a “safe” prediction range (possibly from the pretraining data) when provided with less relevant examples. In the random neighbors setting, the model shows better overall performance and a reduced mode collapse behavior. This suggests that introducing more variability into the training data helps the model to better understand the underlying patterns that predict yield strength. The nearest neighbors setting exhibits the best model performance suggesting that proximity in training data correlates with more accurate yield strength predictions. The mode collapse behavior is significantly reduced compared to the farthest neighbors and random neighbors, reflecting a greater diversity in the model’s output. The observations show varying degrees of mode collapse based on the proximity of prompt examples to the predicted point. For instance, when provided with more closely related training data, the model exhibits stronger predictive signals. This suggests that the extent of mode collapse behavior could serve as an indicator of LLMs’ predictive quality, which can be improved by including more similar examples in the prompt.
Robustness Analysis of Materials Science Q&A
In Figure 6, we present the outcomes of the robustness assessment of gpt-3.5-turbo-0613 when confronted with different types of textual modifications to the MSE-MCQs, to evaluate its stability to various types of "noise". The gpt-3.5-turbo-0613 was chosen because of its moderate performance, availability, and cost-effectiveness.
Figure 6: Robustness analysis of gpt-3.5-turbo-0613 in materials science Q&A using MSE-MCQs dataset. The error bars represent the standard deviation due to LLMs’ non-determinism (Original, Synonym Replacement, Distractive Info) and the randomness (Sentence Reordering, Unit Mixing, Superfluous Info) introduced to the questions.
Ranking by the degradation severity on the easy-level questions, sentence reordering has the least performance drop, followed by synonym replacement, distractive info, unit mixing, and superfluous info. The performance on the hard-level questions is close to the baseline score, indicating the model’s struggle with complex queries regardless of textual modifications, and will not be discussed in detail. The larger error bars in medium and hard questions suggest that LLMs tend to generate more varied responses to complex and lengthy queries.
Sentence reordering has a minimal impact on gpt-3.5-turbo-0613’s performance, demonstrating the model’s robust parsing capabilities. The model maintains comprehension and extracts the necessary information to answer correctly, despite the alteration in the natural flow of questions. This suggests an advanced level of syntactic flexibility and the ability to contextualize information effectively.
The performance drop with synonym replacement indicates that the LLMs can generate different or faulty responses when presented with different terminologies of the same meaning. This reveals LLMs’ reliance on specific terminologies for recognition and comprehension in materials science. In contrast to humans, who can grasp the conceptual continuity behind varied expressions for flexible cognition, LLMs’ struggles with synonym replacement emphasize the need for advanced training that prioritizes semantic networks over mere word recognition [29].
Introducing distractive information mimics a real-world scenario where irrelevant data often accompanies critical information, requiring sharp focus and analytical precision. Improving LLMs’ ability to filter out irrelevant information is crucial for more effective information retrieval, problem-solving, and data interpretation [30]. The results indicate that LLMs struggle to filter distractions in easy-level questions, where mostly simple conceptual knowledge is tested. The added non-essential details seemed to divert the LLMs’ "attention", leading to fewer correct answers. The degradation is less severe in the medium and hard-level questions, where greater complexities and lengths may make the distractive information more distinguishable.
Mixing and converting the units tests LLMs’ abilities to perform numerical reasoning and apply mathematical concepts within a linguistic context. The added complexity introduced by unit mixing degraded the performance of the gpt-3.5-turbo-0613 model on easy and medium-level questions. Although some state-of-the-art LLMs support multi-modal applications and function calls to perform calculations [31], the ability to correctly identify and convert units from a large text can still benefit use cases such as information retrieval and data interpretation.
Superfluous information differs from distractive information in that it is more relevant to the questions themselves. For LLMs, distinguishing the necessary information from merely relevant but non-essential details is a more challenging cognitive process, mirroring the nuances of advanced human problem-solving. It requires an understanding of the problem’s objective, prioritizing information based on the question, and applying only the information that will lead to the correct conclusion. The noted decrease in accuracy shows a current limitation in the LLM’s ability for critical information assessment. The variation in performance degradation is influenced by the type and relevancy of the superfluous information provided.
Robustness Analysis of Materials Property Prediction
Table 1 shows the result of the degradation study on the LLM-Prop model, demonstrating how the LLM’s performance on the band gap prediction is affected by various modifications to the textual descriptions of material crystal structures.
Original Distractive Info Sentence Reordering Superfluous Info
[…]
Table 1: Mean Absolute Error (MAE) under different conditions
After adding distractive information to the material descriptions, the LLM-Prop model showed negligible degradation, indicating this application-specific model can effectively differentiate relevant from irrelevant information. This resilience, likely due to the targeted training and fine-tuning on domain-specific texts, enables it to focus on key features for band gap prediction. This showcases the potential noise-filtering capabilities of the trained and fine-tuned transformer models, which traditional ML models may suffer from.
The impact of sentence reordering increased the MAE by 12.9%, suggesting the model’s reliance on the structured descriptions for accurate predictions. From the previous study on MSE-MCQs degradation, the effect of sentence reordering was less significant, indicating that larger general LLMs, which are trained with more various texts, can exhibit better contextual understanding and are less prone to order changes. The presence of superfluous information, particularly an additional sentence from another material’s description, leads to a 39% increase in MAE. This substantial degradation indicates that while the model can filter out irrelevant distractive noise, it struggles considerably when faced with data that is contextually relevant to the specific prediction task.
To further assess the model’s robustness and determine which description elements are essential for prediction accuracy, we conducted an ablation study that involves altering the structures and lengths of the input description. As shown in Figure 7, the description length is expressed as percent sentence inclusion, ranging from 10% to 100% and MAE is used as a measure of prediction accuracy.
Figure 7: Structure and length effect of LLM-Prop on prediction performance.
When the number of description sentences is incrementally increased, the MAE rapidly decreases as more information is provided, and optimized at 100% sentence inclusion. Interestingly, in the other configurations, the initial MAEs with just 10% sentence inclusion are notably lower and some can double the performance of the original setting. This indicates that the initial sentences may not contain the most useful information for prediction. The MAEs begin to converge around 50% sentence inclusion, beyond which differences become statistically insignificant. Notably, in the random order setting, there is virtually no variation in the MAE when incorporating three different sentence shuffles. This suggests that LLM-Prop can effectively extract key information and deliver consistent predictions, despite variations in the sentence order.
By 40% sentence inclusion, the reversed order yields the lowest MAE, indicating that sentences at the end of descriptions contain crucial information. However, by 50% sentence inclusion, the performance of the reversed order begins to align with that of the random order. This alignment suggests that central information may not be as crucial for prediction accuracy, especially as the random order likely incorporates more of the initial sentences than the reversed order at this inclusion level. Based on these insights, we developed the sides-to-middle approach, aiming to prioritize information at the beginning and the end. The approach shows the most significant improvement, achieving the lowest MAE between 40% and 70% sentence inclusion among all. The error continues to decrease and is optimized at full sentence inclusion being only 5.8% higher than the original setting in MAE. This implies that the contextual framing provided by the beginning and end of the description holds considerable importance for the model’s predictive accuracy, yet the model is still best optimized for the original order.
This ablation study showcases that the fine-tuned model can perform effectively even when provided with significantly reduced prompts. We found that diverging from the training setup (i.e., changing the textual order of the prompt) can sometimes result in improved performance at ablated data volumes. This counterintuitive result suggests that highly templated training or fine-tuning data can lead to unexpected effects. Consequently, this implies two key considerations: 1) training templates should be diverse to prevent models from overfitting to unimportant patterns, and 2) when using a fine-tuned model trained on a specific template, it may not always be optimal to match the template during inference, particularly under ablated data conditions. These insights highlight the potential for optimizing training costs while maintaining performance.
Discussion
The findings of this study offer crucial insights into the behaviors and limitations of LLMs in materials science domain-specific Q&A and materials property prediction. While the robustness analysis indicates that LLMs can manage certain types of noise with resilience, their performance is significantly challenged under more complex and deceptive conditions, such as when superfluous information is introduced. Prompting techniques (e.g. expert prompting, zero-shot chain-of-thought prompting) can sometimes improve the model performance, not by unlocking new capabilities, but by increasing the probability of following an expected format. During property prediction, in-context learning enables pretrained LLMs to achieve relatively high accuracy in low-data settings when provided with few-shot examples with high proximity to the target. However, the observed mode collapse behavior, where the model generates repeated outputs despite varying inputs, showcases that providing ineffective few-shot examples can cause the model to default to a memorized response rather than conditioning its output on the provided prompt. This study also highlights that fine-tuned models can exhibit enhanced performance under ablated data conditions when diverging from the training setup, such as altering the textual order. This underscores the risks associated with fixed templating during fine-tuning and suggests that users should be cautious about strictly matching the training templates during inference. This study highlights the challenges and limitations of using LLMs in materials science, aiming to motivate future improvements that enhance their reliability and effectiveness as research tools.
Methods
The methodology is divided into four subsections that cover the performance evaluation and robustness analysis of LLMs in materials science Q&A and property predictions. In each subsection, we will introduce the models, datasets used, prompting techniques, and evaluation criteria chosen for the specific study.
Performance Evaluation of Materials Science Q&A
Using MSE-MCQs, we benchmarked a range of both commercial and open-source LLMs, including OpenAI’s gpt-3.5-turbo-0613 [32] and gpt-4-0613 [33], alongside Meta AI’s Llama variants - llama2-70b, llama2-13b, and llama2-7b [34]. The suffixes in the Llama model names denote the number of model parameters where, e.g., llama-70b contains 70 billion parameters. This study primarily utilized fixed-version and open-source models to enhance the reproducibility of the findings.
The prompting strategies used in this evaluation are zero-shot chain-of-thought (CoT) and expert prompting. Zero-shot CoT prompting uses an instructional prompt to enable the model to "think aloud" and step through a reasoning process that leads to an answer, potentially boosting the accuracy in problem-solving tasks [35]. Expert prompting, i.e. instructing an LLM to act as a domain expert, has been shown to influence the model’s responses to be more aligned with expert knowledge and reasoning [36]. The specific implementations of each for their respective tasks are defined below.
In the Q&A evaluation, the expert prompt includes instructions to define the domain of study, introduces the settings of the questions, and emphasizes step-by-step reasoning and calculations. The goal is to improve the LLMs’ ability to retrieve domain-specific knowledge, follow the instructions, and correctly perform reasoning and calculations. The expert prompt is shown below:
Expert Prompt: You are a renowned materials science engineering professor with extensive knowledge in the field. Your students have presented you with a challenging multiple-choice question related to materials science engineering. The question requires a detailed understanding and application of materials science principles. Please read the question carefully and provide a step-by-step explanation of your reasoning process, calculations, and analysis. Remember, the question has only one correct answer, which could be option (a), (b), (c), (d), etc. After carefully analyzing and calculating, please present the final answer at the end of your explanation. Your goal is to elucidate the concepts and problem-solving techniques in materials science engineering for your students.
The questions were fed to LLMs in two settings: 1) as is, to benchmark the LLM’s level of understanding of materials science domain-specific knowledge and problem-solving; 2) to evaluate the aforementioned prompting strategies to enhance the LLM performance. Given the lengthy reasoning in the answers and the potential for errors in manual verification, we used the gpt-4-0613 API in a separate client to extract and assess responses automatically. This system compared answers to the provided correct choices, generating a simple binary score (1 for correct, 0 for incorrect) without evaluating the reasoning steps. Finally, the average accuracy of each category was calculated and reported. When selectively compared to manual checks (>200 answers), the method was found to be reliable, consistently identifying correct answers with over 95% accuracy. The prompt is shown below:
System Prompt: You are to read the following text, which is the answer to a multiple choices question. The text should state the final answer (option (a), (b), (c), or (d)). You are to compare the stated answer with the correct answer: ANSWER. If the stated answer is correct, please type 1, otherwise type 0. If the final answer is not one of the options or reports multiple options, it is considered wrong (you should type 0). Do not type anything else other than 1 or 0.
Performance Evaluation of Materials Property Prediction
For this study, we used the matbench_steels dataset to evaluate the predictive capabilities of pretrained LLMs, using their ICL capability with few-shot examples. We benchmarked llama2-7b, llama2-13b, mistral-7b [37], gpt-3.5-turbo-0613, and gpt-4-0613 on their abilities to predict the yield strengths given the steel compositions. For performance comparison, a KNN and an RFR model were also implemented. Few-shot learning involves providing the LLM with a few examples of the task at hand, enabling it to learn the pattern and apply it to unseen questions or problems [38]. To use LLMs as predictive models, we fed the “training data” as few-shot examples to the prompt windows of the LLMs. An example of the “training data” prompt is shown below.
System Prompt: Given some example alloy compositions and their yield strengths, predict the yield strength for one additional alloy composition.
User Prompt: ###Examples###
composition: Fe0.779 Mn0.000101 Si0.000198 Cr0.114 Ni0.0973 Mo0.000116 Nb0.0000597 Co0.0000941 Al0.00432 Ti0.00359, yield strength: <1125.0>
composition: Fe0.673 Mn0.0000992 Si0.0361 Cr0.167 Ni0.0372 Mo0.000114 Nb0.0000587 Co0.0851 Al0.000404 Ti0.000569, yield strength: <1143.7>
Now predict the yield strength for the following composition.
composition: Fe0.752 Mn0.000101 Si0.000197 Cr0.133 Ni0.0783 Mo0.0128 Nb0.0000596 Co0.0000940 Al0.0209 Ti0.00104
Write only the yield strength in a single numerical value that is enclosed by <>
Starting with an instruction, compositions were restructured by separating each element with a space and then paired to their corresponding yield strengths. We varied the number of training data points from 5 to 25 to observe how LLMs’ prediction accuracy scales with data size. Beyond 25 points, some models suffered from limited prompt windows. To assess the impact of data selection variability, five random seeds were chosen to vary the selected data. To compare the predictive capability of LLMs and traditional ML models, we also utilized an RFR model trained using the same data points. The RFR model, unlike the LLMs that operated on raw compositional data, was trained with MAGPIE features [26] extracted from the compositions, aiming to compare the predictive capability of LLMs and traditional ML models. The selected features are presented in the supplementary information. To evaluate the impact of the proximity of the few-shot examples on the predictive performance, a retrieval-augmented method was used. Specifically, the compositions were encoded by their elements and contents and then transformed the compositions into principal components (PCs). Given each test composition, we formulated its “training” set by its Euclidean distances (L2 norm) to the other points in the PC space. Three settings were chosen based on the distances: 1) random; 2) nearest neighbors; 3) farthest neighbors. The prediction accuracy, mean absolute error (MAE), in each setting was computed and analyzed.
Robustness Analysis of Materials Science Q&A
To evaluate the robustness of LLMs for materials science Q&A, we continued using the MSE-MCQs dataset and tested the gpt-3.5-turbo-0613 model to observe how textual perturbations impacted performance. In this study, no prompting strategy was implemented, and the questions were directly used as the user prompts. We identified different types of “noise” that can be introduced to the MCQs to evaluate the robustness of LLMs. As shown in Table 2, the textual inputs were modified systematically in five different ways, i.e. 1) unit mixing, 2) sentence reordering, 3) synonym replacement, 4) distractive info, and 5) superfluous info.
Degradation Type Description Goal
Unit Mixing Mixing and converting the units To test LLMs’ interpretation of different unit systems and calculation abilities
Sentence Reordering Reordering the sentences in the questions To assess LLMs’ capability to maintain comprehension on varied sentence constructions and logical flow
Synonym Replacement Replacing technical nomenclature with their synonyms To evaluate the semantic understanding and stability of LLMs
Distractive Info Adding non-materials-science-related distractive information to the questions To test LLMs’ ability to filter out irrelevant data
Superfluous Info Adding materials-science-related superfluous information containing numerical values to the questions To challenge LLMs’ ability to identify relevant data without being misled by additional numeric details that are potentially incorrect or conflicting
Table 2: Types and descriptions of textual degradation applied to LLMs.
These modifications are expected to vary in their impact on the LLMs’ performance, with some potentially degrading it due to their adversarial nature (such as reordering sentences and adding superfluous materials-science information) and others more realistically simulating conditions encountered in real-life scenarios. Considering the inherent variability due to the non-deterministic nature of LLMs, the test was repeated three times for the original, synonym replacement, and distractive info (same input texts). The unit mixing, sentence reordering, and superfluous info were randomized three times to introduce variability in the data for the evaluation. Finally, the accuracy of each category was calculated and reported.
Robustness Analysis of Materials Property Prediction
In materials property prediction, we selected the LLM-Prop model along with its associated band gap dataset. LLM-Prop is a fine-tuned T5 model, topped with a linear layer, designed to predict materials properties from crystal structure descriptions generated using Robocrystallographer [14, 39, 40].
The material descriptions underwent systematic modifications mirroring those applied in the Q&A evaluations, except for unit mixing and synonym replacement. During data preprocessing for LLM-Prop, all numerical values and units, such as bonding distances and angles, are replaced with a [NUM] token, to emphasize the model’s focus on text-based understanding [14]. Unit mixing might disrupt the preprocessing algorithm, and thus was excluded from the analysis. Synonym replacement was excluded because the original terminology was already highly specific and lacked equivalent synonyms. Furthermore, we conducted an ablation study of textual degradation to examine the model’s resilience against structural and length variations in the input data, as well as to explore which aspects of the descriptions the model relies on for predictions. We manipulated the order and fraction of sentences included, testing configurations including 1) original order, which prioritizes the initial information in a description, 2) reverse order, which prioritizes the sentences from the end of a description, 3) random order, shuffling the information, and 4) sides-to-middle, which deprioritized central information. The impact of these textual degradations was quantitatively assessed by measuring the resultant prediction error in MAE.
Data availability
The data required and generated by our code are publicly available at url (to be inserted upon acceptance of the paper).
Code availability
The code for analysis and figure generations in this work is available on GitHub at url (to be inserted upon acceptance of the paper).
Acknowledgements
We acknowledge Andre Niyongabo Rubungo and Adji Bousso Dieng for providing the dataset and model for the LLM-Prop study.
Author contributions statement
H.W., E.K., and J.H.S. conceived and designed the project. E.K., and J.H.S. supervised the project. H.W. conducted the experiments. S.R. provided the dataset for materials science question & answering study. H.W., K.L., Y.F., E.K., and J.H.S. discussed the results. H.W drafted the manuscript. All authors reviewed and edited the manuscript.
Competing interests
The authors declare no competing interests.